OpenStudy (adamk):

If f(x) = x^2 - 4x + 2, then f(a - b) = ?

OpenStudy (mathmale):

One way of doing this: Write out the function: f(x) = x^2 - 4x + 2 Wherever you see (x) or x, replace it with (b-a). Thus, f(x) becomes f(b-a). Evaluate the right side. Hint: x^2 becomes (b-a)^2. MUST use parentheses.

OpenStudy (adamk):

DUH, I'm should've known that. Brain lapse. Thanks. The a's and b's confused me since I didn't see any 'x'.

OpenStudy (3mar):

Very simple Replace every x with a-b f(a-b)=(a-b)^2-4(a-b)+2 and you can simplify it if you wish

OpenStudy (mathmale):

Typo: I meant a-b, not b-a.

OpenStudy (mathmale):

3mar managed to catch that before I did. Bully for him.